OpenStudy (anonymous):

x^2- sqrt(x)=0 ..... Find the zeroes

OpenStudy (agent0smith):

Add square root of x to both sides, \[\large x^2 = \sqrt x\] But i think you just have to solve it by checking which x values will work here. I can only see 0 and 1.

OpenStudy (agent0smith):

Oh, you can square both sides...\[ \large x^4 = x\] then subtract x on both sides \[\large x^4 - x = 0\] now factor out an x \[\large x (x^3 - 1) =0\] So either x=0 or x^3-1 = 0 Have you done complex numbers?

OpenStudy (agent0smith):

N/m there aren't any complex solutions that will work in the original equation, but i introduced some when i squared both sides.

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):